# TPFL: Tsetlin-Personalized Federated Learning with Confidence-Based Clustering

### v2

@@ Abstract
The world of Machine Learning (ML) has witnessed rapid changes in terms of new models and ways to process users’ data. The majority of work that has been done is focused on Deep Learning (DL) based approaches. However, with the emergence of new algorithms such as the Tsetlin Machine (TM) algorithm, there is growing interest in exploring alternative approaches that may offer unique advantages in certain domains or applications. One of these domains is Federated Learning (FL), in which users’ privacy is of utmost importance. Due to its novelty, FL has seen a surge in the incorporation of personalization techniques to enhance model accuracy while maintaining user privacy under personalized conditions. In this work, we propose a novel approach dubbed TPFL: Tsetlin-Personalized Federated Learning, in which models are grouped into clusters based on their confidence towards a specific class. In this way, clustering can benefit from two key advantages. Firstly, clients share only what they are confident about, resulting in the elimination of wrongful weight aggregation among clients whose data for a specific class may have not been enough during the training. This phenomenon is prevalent when the data are non-Independent and Identically Distributed (non-IID). Secondly, by sharing only weights towards a specific class, communication cost is substantially reduced, making TPLF efficient in terms of both accuracy and communication cost. The results of TPFL demonstrated the highest accuracy on three different datasets; namely MNIST, FashionMNIST and FEMNIST.
@@ 1 Introduction
Today’s world is heavily driven by numerous Machine Learning (ML) algorithms and it is almost impossible to ignore its pervasive existence in our digital lives. This widespread adoption of ML applications equals huge amounts of never-before-seen data from various types of sources, whether it is a face detection camera or a recommendation based on a query in a search engine [1]. This phenomenon in itself raises critical concerns in regards to user privacy as well as data security [2]. Traditional ML techniques rely on centralized approaches, where users data is displaced and moved to where the algorithm resides [3]. The technique suffers from numerous disadvantages, including arduous maintenance of large amounts of data which is produced in mass from various parties, communication cost of data transfer from parties to the centralized server, and finally the most important one, serious privacy risks againsts users data [4].
@@
One of PFL strategies to keep the personalized preferences of clients intact during the aggregation is multi-center FL, in which similar clients are partitioned into clusters. Thus, unlike single-center FL where there is only one single global model for all the clients, with multi-center FL we can take personalization into account and group similar clients into their corresponding cluster. This multi-center partitioning establishes numerous clusters, each is responsible for weight aggregation of its corresponding clients.
Numerous clustering based personalization strategies have been proposed. Ghosh et al. in [11] introduce a one-shot clustering algorithm called Iterative Federated Clustering Algorithm (IFCA) to address the challenges in FL clustering using Empirical Risk Minimizers (ERM) from FL clients. They also discuss practical aspects of implementing IFCA, including a weight-sharing scheme that reduces communication costs between the aggregator and FL clients. The IFCA uses FEMNIST and MNIST for its work, which yields satisfactory accuracy in all experimental setups. Work of Briggs et al. introduces a novel method for training specialized ML models over distributed datasets enhanced by hierarchical clustering [12]. Their paper continues to add a clustering step within the federated learning protocol, which was clustering clients based on the similarity of their model weight updates. Another positive aspect of this work was comprehensive analysis of how hierarchical clustering impacts test set accuracy across different iid and non-iid data settings. Another valuable work for clustering-based PFL was the work of Ruan and Joe-Wong in [13], which proposed FedSoft. Their contribution was based on proximal local updates. Unlike traditional hard clustering, where each client belongs to a single cluster, FedSoft allows clients to be associated with multiple clusters with varying degrees of membership. A proximal term is added to the local objective function of each client, ensuring the local models stay close to the global model, which as a result helps in mitigating the divergence issues often encountered in federated learning due to heterogeneous data distributions across clients.
In this paper, we present a novel FL method that, to the best of our knowledge, leverages the Tsetlin Machine (TM) algorithm for the first time in the context of personalization, which we dub Tsetlin-Personalized Federated Learning (TPFL). This innovative approach is designed to perform effectively with both iid and non-iid data in a few-shot learning fashion. In essence, the clustering mechanism of TPFL is directly based on each client’s confidence towards a specific class of data. In fact, clients generally in an FL setting tend to be trained on data that they generate locally. Therefore, a client that is trained on images of odd numbers tends to have more confidence towards those classes of data. And similarly a client with images of even numbers is confident about what it had been trained on; a range of even numbers. As a result, the more data the local model is trained on, the more confident it becomes towards that particular class. Our novelty specially holds true when FL clients data distribution is non-iid. In other words, when some random clients are only trained on data whose distribution is heavily skewed towards a similar class of data, they must be confident and consequently partitioned into the same cluster.
@@
  • We formulate an approach to measure the confidence of each client after they have been trained.
  • In the proposed method, we use maximization, where the model’s highest confidence towards a specific class determines the cluster to which the client belongs.
  • Clients are partitioned into clusters based on their confidence towards specific labels of the data. Therefore, the total number of clusters equals the number of data classes.
  • We evaluate the TPFL method under 5 different experimental setups on 3 datasets, allowing us to see the effectiveness of TPFL.
  • We provide an algorithm for both the aggregator and the clients to complete the training in 10 rounds only. This allows us to see the efficiency of the TPFL compared to baseline models.
The organization of this paper is as follows: section 2 provides a review on related works that proposed a novel method for PFL, secion 3 discusses the motivation behind this work, section 4 goes over the preliminaries, section 5 provides a comprehensive explanation on the proposed TPFL method and section 6 provides the obtained results of our implementation of TPFL and other baseline methods.
@@ 2 Related Works
Since the advent of FL-based architecture, different approaches have opted for PFL to increase the efficiency and accuracy of FL models. The work of Long et al. brings forward an approach called Federated Stochastic Expectation Maximization (FeSEM) that solves the challenge of non-IID data using a framework with multi-center clusters for clients [14]. This approach partitions clients based on their data similarity using K-Means algorithm that is guided by the FeSEM algorithm. Effectively, each user is aligned with its corresponding global model and as a result, the loss is minimized in each cluster. The authors used Federated Extended MNIST (FEMNIST), FedCelebA datasets for their work and case study comparison to illustrate the effectiveness of their results. The authors however, did not show the communication cost of their model since one of the important criteria within FL settings is the communication cost between the clients and the aggregator, whether the FL setting is based on a single-center or a multi-center architecture.
@@
Work of Mansour et al. [15] proposed three distinct approaches to achieve better personalization in FL settings which are user clustering, data interpolation and model interpolation, respectively. Although the approaches are not combined, their comparison sheds a light on different aspects of personalization. The clustering scheme allows the FL architecture to be more scalable than the predecessor approaches. Their benchmark datasets were EMNIST-62 as well as synthetic data that was used for the initial evaluation and understanding of the algorithms’ behavior. Sattler et al. proposed a novel model-agnostic clustering algorithm for clients whose distributions are similar when it comes to non-iid data [16]. Their unique approach is the adaptiveness of their framework, which adjusts the clusters as more data becomes available. They used MNIST, CIFAR-10 and Fashion-MNIST (FMNIST) datasets for their framework. Gong et al. proposed another framework for clustering clients based on non-iid data [17]. Their novel strategies include a local training adjustment strategy as well as an adaptive weighted strategy for client clustering based on voting. Model accuracy was improved and communication overhead was reduced by evaluating the model on MNIST, FMNIST and CIFAR-10 datasets. Morafah et al. have proposed yet another federated framework for non-iid data called Federated Learning by Inference Similarity (FLIS) [18]. Their clustering approach is based on inference similarity. Their approach does not force the number of clusters to be known prior to the cluster creation process. Their model-agnostic approach minimizes the need for frequent communication between clients and the central server. Their work focuses on MNIST, FMNIST and CIFAR-10. Cho et al. have proposed a framework called COMET that takes the local models’ data as well as their architecture into consideration, which is likely to be a scenario for industrial FL architecture [19]. Moreover, their approach relied on knowledge distillation mechanism that allows clients to share only the logits (predictions) on a common unlabeled dataset. In addition to reduced communication overhead, their algorithm’s improved accuracy outperforms some of the contenders in FL settings. Their work was trained on CIFAR-10 and CIFAR-100 datasets. Work of Fraboni et al. aims to enhance client representativity and reduce the variance in the stochastic aggregation weights of clients. This framework is based on 1) sample size-based clustering, which groups clients based on the size of their datasets and 2) model similarity-based clustering, which groups clients based on the similarity of their model updates. The framework is trained on MNIST and CIFAR-10 datasets [20]. Work of Camara et al. proposes a clustered FL architecture specifically designed for unsupervised anomaly detection in large, heterogeneous networks of IoT devices. This approach addresses the challenges of data isolation, network overhead, and privacy concerns. Their results demonstrate the improvements in anomaly detection performance in addition to reduction of network overhead and improved global model convergence. Their work was evaluated on a benign IoT network traffic testbed called Gotham [21]. Work of Yin et al. introduced an efficient one-off clustered federated learning framework called FedEOC, which aims to improve the balance between accuracy and efficiency in federated learning by clustering clients only once and applying the clusters over multiple rounds. FedEOC leverages the learning-to-learn characteristic of meta-learning in order to quickly adapt to new tasks, ensuring that the initial model parameters are well-suited for all clients. The FedEOC framework was evaluated on the ml-100K and ml-1M dataset [22]. Li et al. put forward a new algorithm called Federated Learning with Soft Clustering (FLSC). This algorithm combines the strengths of soft clustering and the Iterative Federated Clustering Algorithm (IFCA). Contrary to IFCA, which uses non overlapping clusters, FLSC allows for overlapping clusters. This means that each client can contribute to multiple clusters, improving the utilization of local information [23]. Their work used MNIST and FMNIST datasets. He et al. proposed a framework that brings forward a novel layer-wise perturbation function that reduces the Local Differential Privacy (LDP) noise in comparison to relevant methods. As a result, communication overhead was decreased in Internet of Things (IoT) devices. MNIST, fashion-MNIST (FMNIST) and Federated Extended MNIST (FEMNIST) were used for the evaluation of their framework [24]. Finally, the work of Ghosh et al. addresses the challenges of robustness in a heterogeneous environment with Byzantine nodes. Their proposed method consists of a three-stage modular algorithm; 1) Worker nodes compute their Local Empirical Risk minimizers (ERMs) and send them to the center node, 2) The center node clusters these local ERMs to form distinct clusters and 3) Each cluster undergoes a Byzantine-tolerant distributed optimization process to refine the model parameters. Results with 3 different clustering strategies were presented, namely K-means (KM), Trimmed K-means (TKM), and K-geomedians (KGM). The results demonstrate that the proposed algorithm outperforms standard non-robust algorithms in both synthetic and real-world datasets. The real-world dataset that was used was Yahoo! Learning to Rank Dataset [25].
@@ 3 Motivation
The advent of Federated Learning (FL) has paved the way to the collaboration of training among multiple Artificial Intelligence (AI) models while maintaining user data privacy. Recent advancements within FL have led many to propose personalization techniques in order to increase the efficiency of FL models. Nevertheless, the potential use of other state-of-the-art algorithms like the TM algorithm is normally overlooked by the widespread use of Deep Learning (DL) models within the FL context.
@@ 4.1 Tsetlin Machine
in which $c_{j}^{+}$ is the vote summation of all positive-polarity clauses and $c_{j}^{-}$ is the vote summation of all negative-polarity clauses. Also, $w_{j}^{+}$ is the weight of the $j^{th}$ clause with positive polarity and $w_{j}^{-}$ is the weight of the $j^{th}$ clause with negative polarity [30].
@@ 4.2 Problem Statement
Each client $i$ is assigned to a cluster $k$ where $k\in\{1,2,\ldots,C\}$ based on the class $l$ for which they show the highest confidence score. Client $i$ is assigned to cluster $k$ such that:
$k=\arg\max_{1\leq l\leq C}\text{conf}_{i,l}$
The final outcome of this clustering method is $C$ clusters, where each cluster $k$ contains clients whose highest confidence score corresponds to class $k$. For each cluster $k$, the clients weights within the cluster are aggregated and averaged. We denote the model weights of client $i$ as $\theta_{i}$. The aggregated models weights for cluster $k$ are denoted as:
$\theta_{k}=\frac{1}{|K_{k}|}\sum_{i\in K_{k}}\theta_{i}$
@@ 5.1 Phase A
Phase A is where the process occurs on the local device. Phase A consists of 5 steps in total. In phase A, in the first step, each client starts training its local model using its private data.
@@
The client $m$ initiate the positive-polarity clauses $C_{i}^{+}$ and the negative-polarity clauses $C_{i}^{-}$ as well as its weights $W_{m}[c]$ for each class $c$. The TM clauses are updated during each round using literals from the local training set $D_{\text{train}}$. In step 2 within each iteration, the client updates the clause weights for each class $c$ based on local data $D_{\text{train}}$ and calculates a predicted output $\hat{y}$. This process continues over several desired epochs to improve the classification accuracy of the local model. In step 3, the confidence score is locally calculated using a separate set of data called $D_{\text{conf}}$. This confidence stems from the voting machanism of TM algorithm. The more votes there are towards a class $c$, the more confident the model is in its prediction $\hat{y}$. Afterwards, In step 4 the local model selects the maximum confidence and sends a tuple to the aggregator, containing the weight vector $W_{m}^{c}$ and its corresponding class $c$ to the aggregator. In this way each client sends only the weight vector $W_{m}^{c}$ associated with the class $c$ for which it has the highest confidence. This selective transmission optimizes communication efficiency while ensuring that each client contributes meaningfully to the entire cluster in areas where it has the strongest expertise. Algorithm 1 demonstrates phase A in the local client.
@@
Require: Number of clients $m$, number of classes $c$, number of clusters $k$, weight vector $W_{m}[c]$, positive-polarity clauses $C_{i}^{+}$, negaive-polarity clauses $C_{i}^{-}$, local predicted class $\hat{y}[m]$.
1:Input: Local dataset $D_{\text{train}}$, confidence set $D_{\text{conf}}$.
2:Output: Updated weight vector $W_{m}^{{c}_{\text{max}}}$ for class $c_{\text{max}}$ with highest confidence.
3:Initialize clauses positive-polarity clauses $C_{i}^{+}$ and negaive-polarity clauses $C_{i}^{-}$ and weights $W_{m}[c]$ for each client $m$ and each class $c$.
4:for local training epoch do
5:     Update clause weights $W_{m}[c]$ based on $D_{\text{train}}$.
6:     Compute local prediction:
$\hat{y}[m]=\arg\max_{c}\left(\sum_{j=1}^{n/2}W_{j}^{i,+}C_{j}^{i,+}-\sum_{j=1}^{n/2}W_{j}^{i,-}C_{j}^{i,-}\right).$
7:end for
8:Compute confidence score for each class $c$ on $D_{\text{conf}}$:
$c_{\text{max}}=\max_{1<c<C}\left(\sum_{x\in D_{\text{conf}}}\left(\sum_{j=1}^{n/2}C_{j}^{i,+}(x)-\sum_{j=1}^{n/2}C_{j}^{i,-}(x)\right)\right).$
9:Select class $c$ with highest confidence score $c_{\text{max}}$.
10:Send weight vector $W_{m}^{{c}_{\text{max}}}$ for class $c_{\text{max}}$ to aggregator.

### v3

@@ Abstract
The world of Machine Learning (ML) has witnessed rapid changes in terms of new models and ways to process users’ data. The majority of work that has been done is focused on Deep Learning (DL) based approaches. However, with the emergence of new algorithms such as the Tsetlin Machine (TM) algorithm, there is growing interest in exploring alternative approaches that may offer unique advantages in certain domains or applications. One of these domains is Federated Learning (FL), in which users’ privacy is of utmost importance. Due to its novelty, FL has seen a surge in the incorporation of personalization techniques to enhance model accuracy while maintaining user privacy under personalized conditions. In this work, we propose a novel approach called TPFL: Tsetlin-Personalized Federated Learning, in which models are grouped into clusters based on their confidence towards a specific class. In this way, clustering can benefit from two key advantages. Firstly, clients share only what they are confident about, resulting in the elimination of wrongful weight aggregation among clients whose data for a specific class may have not been enough during the training. This phenomenon is prevalent when the data are non-Independent and Identically Distributed (non-IID). Secondly, by sharing only weights towards a specific class, communication cost is substantially reduced, making TPLF efficient in terms of both accuracy and communication cost. The TPFL results were compared with 6 other baseline methods; namely FedAvg, FedProx, FLIS DC, FLIS HC, IFCA and FedTM. The results demonstrated that TPFL performance better than baseline methods with 98.94% accuracy on MNIST, 98.52% accuracy on FashionMNIST and 91.16% accuracy on FEMNIST dataset.
@@ 1 Introduction
Today’s world is heavily driven by numerous Machine Learning (ML) algorithms and it is almost impossible to ignore its pervasive existence in our digital lives. This widespread adoption of ML applications equals huge amounts of never-before-seen data from various types of sources, whether it is a face detection camera or a recommendation based on a query in a search engine w1. This phenomenon in itself raises critical concerns in regards to user privacy as well as data security [2]. Traditional ML techniques rely on centralized approaches, where users data is displaced and moved to where the algorithm resides [3]. The technique suffers from numerous disadvantages, including arduous maintenance of large amounts of data which is produced in mass from various parties, communication cost of data transfer from parties to the centralized server, and finally the most important one, serious privacy risks againsts users data [4].
@@
One of PFL strategies to keep the personalized preferences of clients intact during the aggregation is multi-center FL, in which similar clients are partitioned into clusters. Thus, unlike single-center FL where there is only one single global model for all the clients, with multi-center FL we can take personalization into account and group similar clients into their corresponding cluster. This multi-center partitioning establishes numerous clusters, each responsible for weight aggregation of its corresponding clients. One of the best ML algorithms that can be used in FL settings is the Tsetlin Machine (TM). The TM algorithm is a great choice compared to Deep Learning (DL) when it comes to interpretability and explainability. This phenomenon stems from the fact that the TM relies on a voting mechanism that enables the TM models to be more transparent in comparison with DL models that possess a black-box nature [11]. In this paper, we present a novel multi-center FL method that, to the best of our knowledge, leverages the Tsetlin Machine (TM) algorithm for the first time in the context of PFL, which we call Tsetlin-Personalized Federated Learning (TPFL). This innovative approach is designed to perform effectively with both IID and non-IID data in a few-shot learning fashion. In essence, the clustering mechanism of TPFL is directly based on each client’s confidence towards a specific class of data. In fact, clients generally in an FL setting tend to be trained on data that they generate locally. Therefore, a client that is trained on images of odd numbers tends to have more confidence towards those classes of data. And similarly a client with images of even numbers is confident about what it had been trained on; a range of even numbers. As a result, the more data the local model is trained on, the more confident it becomes towards that particular class. Our novelty specially holds true when FL clients data distribution is non-IID. In other words, when some random clients are only trained on data whose distribution is heavily skewed towards a similar class of data, they must be confident and consequently partitioned into the same cluster.
@@
  • We provide a client-side algorithm that shows how a client participates in each round of training by sharing only a single weight vector per round. This vector corresponds to the class towards which the client’s confidence is highest.
  • We also provide an algorithm in the aggregator side for the aggregation of clients weights. This indicates that the most number of clusters that we may have is equal to the number of classes.
  • We evaluate the TPFL method under 5 different experimental setups on 3 datasets, showing the effectiveness of TPFL.
  • We compare the TPFL result with 6 other baseline methods. This comparison covers not only the methods accuracy but also their communication costs as well .
The organization of this paper is as follows: Section 2 provides a review on related works that proposed a novel method for PFL, Section 3 discusses the motivation behind this work, Section 4 goes over the preliminaries, Section 5 provides a comprehensive explanation on the proposed TPFL method and Section 6 provides the obtained results of our implementation of TPFL and other baseline methods.
@@ 2 Related Works
Since the advent of FL-based architecture, different approaches have opted for PFL to increase the efficiency and accuracy of FL models. Ghosh et al. in [12] introduce a one-shot clustering algorithm called Iterative Federated Clustering Algorithm (IFCA) to address the challenges in FL clustering using Empirical Risk Minimizers (ERM) from FL clients. They also discuss practical aspects of implementing IFCA, including a weight-sharing scheme that reduces communication costs between the aggregator and FL clients. The IFCA uses FEMNIST and MNIST for its work, which yields satisfactory accuracy in all experimental setups. Work of Briggs et al. introduces a novel method for training specialized ML models over distributed datasets enhanced by hierarchical clustering [13]. Their paper continues to add a clustering step within the federated learning protocol, which was clustering clients based on the similarity of their model weight updates. Another positive aspect of this work was comprehensive analysis of how hierarchical clustering impacts test set accuracy across different IID and non-IID data settings. Another valuable work for clustering-based PFL was the work of Ruan and Joe-Wong in [14], which proposed FedSoft. Their contribution was based on proximal local updates. Unlike traditional hard clustering, where each client belongs to a single cluster, FedSoft allows clients to be associated with multiple clusters with varying degrees of membership. A proximal term is added to the local objective function of each client, ensuring the local models stay close to the global model, which as a result helps in mitigating the divergence issues often encountered in federated learning due to heterogeneous data distributions across clients.
@@
The work of Long et al. brings forward an approach called Federated Stochastic Expectation Maximization (FeSEM) that solves the challenge of non-IID data using a framework with multi-center clusters for clients [15]. This approach partitions clients based on their data similarity using K-Means algorithm that is guided by the FeSEM algorithm. Effectively, each user is aligned with its corresponding global model and as a result, the loss is minimized in each cluster. The authors used Federated Extended MNIST (FEMNIST), FedCelebA datasets for their work and case study comparison to illustrate the effectiveness of their results. The authors however, did not show the communication cost of their model since one of the important criteria within FL settings is the communication cost between the clients and the aggregator, whether the FL setting is based on a single-center or a multi-center architecture. Work of Mansour et al. [16] proposed three distinct approaches to achieve better personalization in FL settings which are user clustering, data interpolation and model interpolation, respectively. Although the approaches are not combined, their comparison sheds a light on different aspects of personalization. The clustering scheme allows the FL architecture to be more scalable than the predecessor approaches. Their benchmark datasets were EMNIST-62 as well as synthetic data that was used for the initial evaluation and understanding of the algorithms’ behavior. Sattler et al. proposed a novel model-agnostic clustering algorithm for clients whose distributions are similar when it comes to non-IID data [17]. Their unique approach is the adaptiveness of their framework, which adjusts the clusters as more data becomes available. They used MNIST, CIFAR-10 and Fashion-MNIST (FMNIST) datasets for their framework. Gong et al. proposed another framework for clustering clients based on non-IID data [18]. Their novel strategies include a local training adjustment strategy as well as an adaptive weighted strategy for client clustering based on voting. Model accuracy was improved and communication overhead was reduced by evaluating the model on MNIST, FMNIST and CIFAR-10 datasets.
@@
Morafah et al. have proposed yet another federated framework for non-IID data called Federated Learning by Inference Similarity (FLIS) [19]. Their clustering approach is based on inference similarity. Their approach does not force the number of clusters to be known prior to the cluster creation process. Their model-agnostic approach minimizes the need for frequent communication between clients and the central server. Their work focuses on MNIST, FMNIST and CIFAR-10. Cho et al. have proposed a framework called COMET that takes the local models’ data as well as their architecture into consideration, which is likely to be a scenario for industrial FL architecture [20]. Moreover, their approach relied on knowledge distillation mechanism that allows clients to share only the logits (predictions) on a common unlabeled dataset. In addition to reduced communication overhead, their algorithm’s improved accuracy outperforms some of the contenders in FL settings. Their work was trained on CIFAR-10 and CIFAR-100 datasets. Work of Fraboni et al. aims to enhance client representativity and reduce the variance in the stochastic aggregation weights of clients. This framework is based on 1) sample size-based clustering, which groups clients based on the size of their datasets and 2) model similarity-based clustering, which groups clients based on the similarity of their model updates. The framework is trained on MNIST and CIFAR-10 datasets [21]. Work of Camara et al. proposes a clustered FL architecture specifically designed for unsupervised anomaly detection in large, heterogeneous networks of IoT devices. This approach addresses the challenges of data isolation, network overhead, and privacy concerns. Their results demonstrate the improvements in anomaly detection performance in addition to reduction of network overhead and improved global model convergence. Their work was evaluated on a benign IoT network traffic testbed called Gotham [22]. Work of Yin et al. introduced an efficient one-off clustered federated learning framework called FedEOC, which aims to improve the balance between accuracy and efficiency in federated learning by clustering clients only once and applying the clusters over multiple rounds. FedEOC leverages the learning-to-learn characteristic of meta-learning in order to quickly adapt to new tasks, ensuring that the initial model parameters are well-suited for all clients. The FedEOC framework was evaluated on the ml-100K and ml-1M dataset [23]. Li et al. put forward a new algorithm called Federated Learning with Soft Clustering (FLSC). This algorithm combines the strengths of soft clustering and the Iterative Federated Clustering Algorithm (IFCA). Contrary to IFCA, which uses non overlapping clusters, FLSC allows for overlapping clusters. This means that each client can contribute to multiple clusters, improving the utilization of local information [24]. Their work used MNIST and FMNIST datasets. He et al. proposed a framework that brings forward a novel layer-wise perturbation function that reduces the Local Differential Privacy (LDP) noise in comparison to relevant methods. As a result, communication overhead was decreased in Internet of Things (IoT) devices. MNIST, fashion-MNIST (FMNIST) and Federated Extended MNIST (FEMNIST) were used for the evaluation of their framework [25]. Finally, the work of Ghosh et al. addresses the challenges of robustness in a heterogeneous environment with Byzantine nodes. Their proposed method consists of a three-stage modular algorithm; 1) Worker nodes compute their Local Empirical Risk minimizers (ERMs) and send them to the center node, 2) The center node clusters these local ERMs to form distinct clusters and 3) Each cluster undergoes a Byzantine-tolerant distributed optimization process to refine the model parameters. Results with 3 different clustering strategies were presented, namely K-means (KM), Trimmed K-means (TKM), and K-geomedians (KGM). The results demonstrate that the proposed algorithm outperforms standard non-robust algorithms in both synthetic and real-world datasets. The real-world dataset that was used was Yahoo! Learning to Rank Dataset [26].
@@ 3 Motivation
The advent of Federated Learning (FL) has paved the way to the collaboration of training among multiple Artificial Intelligence (AI) models while maintaining user data privacy. Recent advancements within FL have led many to propose personalization techniques in order to increase the efficiency of FL models. Nevertheless, the potential use of other state-of-the-art algorithms like the TM algorithm is normally overlooked by the widespread use of DL models within the FL context.
@@ 4.1 Tsetlin Machine
in which $c_{j}^{+}$ is the vote summation of all positive-polarity clauses for class $i$ and $c_{j}^{-}$ is the vote summation of all negative-polarity clauses for class $i$. Also, $w_{j}^{+}$ is the weight of the $j^{th}$ clause with positive polarity for class $i$ and $w_{j}^{-}$ is the weight of the $j^{th}$ clause with negative polarity for class $i$ [31].
@@ 4.2 Problem Statement
Each client $i$ is assigned to a cluster $k$ where $k\in\{1,2,\ldots,C\}$ for which they show the highest confidence score. Client $i$ is assigned to cluster $k$ such that:
$k=\arg\max_{1\leq l\leq C}\text{conf}_{i,l},$
The final outcome of this clustering method is at most $C$ clusters, where each cluster $k$ contains clients whose highest confidence score corresponds to class $k$. For each cluster $k$, the clients weights within the cluster are aggregated. We denote the model weights of client $i$ as $\theta_{i}$. The aggregated models weights for cluster $k$ are denoted as:
$\theta_{k}=\frac{1}{|K_{k}|}\sum_{i\in K_{k}}\theta_{i},$
@@ 5.1 Phase A
Phase A is where the process occurs on the local device. Phase A consists of 5 steps in total. In Phase A, in Step 1, each client starts training its local model using its private data.
@@
The client $m$ initiates the positive-polarity clauses $C_{j}^{+}$ and the negative-polarity clauses $C_{j}^{-}$ as well as its weights $W_{m}[c]$ for each class $c$. The TM clauses are updated during each round using literals from the local training set $D_{\text{train}}$. In Step 2 within each iteration, the client updates the clause weights for each class $c$ based on local data $D_{\text{train}}$ and calculates a predicted output $\hat{y}$. This process continues over several desired epochs to improve the classification accuracy of the local model. In Step 3, the confidence score is locally calculated using a separate set of data called $D_{\text{conf}}$. This confidence stems from the voting machanism of TM algorithm. The more votes there are towards a class $c$, the more confident the model is in its prediction $\hat{y}$. Afterwards, In Step 4 the local model selects the maximum confidence and sends a tuple to the aggregator, containing the weight vector $W_{m}^{c}$ and its corresponding class $c$ to the aggregator. In this way each client sends only the weight vector $W_{m}^{c}$ associated with the class $c$ for which it has the highest confidence. This selective transmission optimizes communication efficiency while ensuring that each client contributes meaningfully to the entire cluster in areas where it has the strongest expertise. Algorithm 1 demonstrates Phase A in the local client.
@@
Require: Number of clients $m$, number of classes $C$, weight vector $W_{m}[c]$, positive-polarity clauses $C_{j}^{+}$, negative-polarity clauses $C_{j}^{-}$, local predicted class $\hat{y}[m]$.
Input: Local train set $D_{\text{train}}$, confidence set $D_{\text{conf}}$.
Output: Updated weight vector $W_{m}^{{c}_{\text{max}}}$ for class $c_{\text{max}}$ with highest confidence.
1:Initialize $C_{j}^{+}$, $C_{j}^{-}$, and $W_{m}[c]$ for each class $c\in\{1,2,\ldots,C\}$.
2:for local training epoch do
3:     Update clause weights $W_{m}[c]$ based on $D_{\text{train}}$.
4:     Compute local prediction:
$\hat{y}[m]=\arg\max_{1\leq c\leq C}\left(\sum_{j=1}^{n/2}W_{j}^{c,+}C_{j}^{c,+}-\sum_{j=1}^{n/2}W_{j}^{c,-}C_{j}^{c,-}\right).$
5:end for
6:Compute confidence score for each class $c$ on $D_{\text{conf}}$:
$c_{\text{max}}=\arg\max_{1\leq c\leq C}\left(\sum_{x\in D_{\text{conf}}}\left(\sum_{j=1}^{n/2}C_{j}^{c,+}(x)-\sum_{j=1}^{n/2}C_{j}^{c,-}(x)\right)\right).$
7:Send weight vector $W_{m}^{{c}_{\text{max}}}$ for class $c_{\text{max}}$ to aggregator.